# Boosted GFlowNets Improving Exploration via Sequential Learning

### v1

@@ 1 Introduction
Generative Flow Networks (GFlowNets) learn stochastic policies over directed acyclic state graphs to sample structured objects in proportion to unnormalized terminal rewards (bengio2021). This framework makes it possible to discover diverse candidates even when feedback is only available at the end of a trajectory, and has shown promise in scientific discovery and combinatorial design. In practice, however, training on large, multimodal targets is often bottlenecked by exploration: the learner rapidly concentrates on easy-to-reach modes while leaving hard-to-reach ones severely undercovered, limiting both coverage and generalization.
@@
Contributions. Our main contributions are as follows:
@@ 2 Related Work
GFlowNets face well-documented exploration bottlenecks on large, multimodal targets. Practical remedies span three main angles: (i) auxiliary exploration policies, e.g., sibling-augmented schemes that collect novelty-seeking trajectories, often via intrinsic rewards such as Random Network Distillation, and re-label them off-policy for training the main learner (madan2025-sa-gfn); (ii) adaptive curricula, where teacher–student mechanisms steer sampling toward under-covered regions identified by high loss or discrepancy signals (kim2025-adaptive-teacher); and (iii) loss-design perspectives, where alternative regression losses (zero-avoiding families) are motivated through divergence analyses to bias training away from mode collapse and toward broader coverage (hu2025-beyond-squared). Scaling-oriented approaches decompose the state graph and train subgraphs asynchronously to enlarge visited regions under fixed budgets (silva2025generalization). Complementary theoretical works study conditions for correctness and distributional mismatch in finite-budget regimes, clarifying when coverage failures arise in practice (silva2025when).
Boosting Variational Inference (BVI; guo2016boosting; miller17boosting; locatello18boosting) generalizes classical boosting to probabilistic inference by iteratively constructing a mixture approximation to the target posterior. At each round, a new variational component is added to reduce the overall Kullback–Leibler divergence between the mixture and the true distribution. This procedure can be interpreted as performing functional gradient descent in the space of probability measures, where each component incrementally corrects the residual approximation error of the current mixture. Under mild smoothness and compactness assumptions on the variational family, BVI enjoys explicit convergence guarantees and systematically improves multimodal coverage by expanding the support of the approximation.
@@ 3 Background and Preliminaries
Generative Flow Networks (GFlowNets, bengio2021) are probabilistic models designed to sample discrete objects $x\in\mathcal{X}$ with probability proportional to a non-negative reward function $R(x)$. Unlike standard generative models, which typically learn a normalized distribution, GFlowNets operate by training a pair of forward and backward policies over trajectories $\tau=(s_{0}\to s_{1}\to\dots\to s_{T})$ in a directed acyclic graph (DAG) of states. The forward policy $P_{F}$ induces a distribution over terminal states, while the backward policy $P_{B}$ provides a stochastic inverse, together defining a system of “flows” that approximate $R(x)$. This formulation has been shown effective in domains where rewards are sparse, multimodal, or expensive to evaluate, making GFlowNets a promising tool for scientific discovery and structured generative modeling.
Formally, let $\Gamma$ denote the set of all valid trajectories starting from the initial state $s_{0}$ and ending in any terminal state. Each trajectory $\tau\in\Gamma$ is assigned a forward probability $P_{F}(\tau)=\prod_{t=0}^{T-1}P_{F}(s_{t+1}\mid s_{t})$ and a corresponding backward probability $P_{B}(\tau\mid x)=\prod_{t=0}^{T-1}P_{B}(s_{t}\mid s_{t+1})$, where $s_{T}=x$. The marginal probability $P_{F}(x)$ of sampling a terminal state $x$ is given by the sum of probabilities of all trajectories $\tau$ that end in $x$:
$P_{F}(x)=\sum_{\tau\in\Gamma:s_{T}(\tau)=x}P_{F}(\tau),$ (1)
where $s_{T}(\tau)$ denotes the terminal state of trajectory $\tau$. A GFlowNet is trained so that this marginal probability satisfies
@@
with $Z=\sum_{x\in\mathcal{X}}R(x)$ acting as the normalizing constant. The objective of training is therefore to align the induced terminal distribution $P_{F}$ with the unnormalized target defined by $R(x)$.
Training objectives. Several training objectives have been proposed to enforce the proportionality between $P_{F}(x)$ and $R(x)$, including Detailed Balance (DB, malkin2022trajectory), Subtrajectory Balance (SubTB, madan2023learninggflownetspartialepisodes). Our work builds on the widely used Trajectory Balance condition (TB, malkin2022trajectory). The TB objective introduces a scalar parameter $Z_{\theta}$ and enforces the following balance condition along full trajectories:
@@ 4.2 Induced Reward and Residual Principle
Let $\Gamma_{x}$ denote the set of trajectories that terminate at $x$. For a trained (frozen) stage with parameters $\theta$, define the per-trajectory estimator
$\widehat{R}_{\theta}(x;\tau)=Z_{\theta}\,\frac{P_{F}^{\theta}(\tau)}{P_{B}^{\theta}(\tau\mid x)},\quad\tau\in\Gamma_{x}.$ (7)
@@
$\widehat{R}_{\theta}(x):=\mathbb{E}_{\tau\sim P_{B}^{\theta}(\cdot\mid x)}\big[\widehat{R}_{\theta}(x;\tau)\big].$ (8)
@@
Let $S:=\{x\in\mathcal{X}:P_{F}(x)>0\}$ be the support of the forward policy’s terminal distribution. Assume that the Trajectory Balance loss has zero expectation, $\mathbb{E}_{\tau\sim P_{F}}[\mathcal{L}_{TB}(\tau)]=0$. Further assume that for every terminal $x\in S$, the backward policy $P_{B}(\cdot|x)$ and the restriction of the forward policy $P_{F}(\cdot)|_{\Gamma_{x}}$ are mutually absolutely continuous.Then, for every terminal $x\in\mathcal{X}$:
$\mathbb{E}_{\tau\sim P_{B}(\cdot|x)}[\widehat{R}_{\theta}(x;\tau)]=R(x)\cdot\mathbb{I}[x\in S]$
and,
$\operatorname{Var}_{\tau\sim P_{B}(\cdot|x)}[\widehat{R}_{\theta}(x;\tau)]=0$
@@
This yields the residual principle: given a frozen stage (or ensemble) that already matches parts of the target, its induced estimate, which we denote $\hat{R}_{old}(x)$, can be used to define a residual target. The next stage is then trained to capture this residual, thereby reallocating learning pressure toward under-covered regions while preserving coverage in regions that are already matched. This principle is operationalized in our boosted loss, detailed next.
@@ 4.3 Boosted Trajectory Balance Loss
Our sequential training framework is motivated by a simple observation: the overall objective is to enforce the Trajectory Balance (TB) condition (Eq. (3)) over the set of all possible trajectories.
When a GFlowNet minimizes the expected Trajectory Balance (TB) loss, $\mathbb{E}_{\tau\sim P_{\mathrm{F}}}[\mathcal{L}_{\text{TB}}(\tau)]$ to zero the TB condition (Eq. (3)) is satisfied for all trajectories within the support of the forward policy, $P_{\mathrm{F}}$. Let $S_{\text{old}}$ be the set of terminal states covered by a converged past GFlowNet, and let $\Gamma_{S_{\text{old}}}\subset\Gamma$ be the set of trajectories ending in $S_{\text{old}}$. The remaining learning objective for a new GFlowNet is thus to enforce the same balance condition on the complement set of trajectories, $\Gamma\setminus\Gamma_{S_{\text{old}}}$, which lead to the under-covered modes.
@@
$R(x)=\underbrace{R(x)\cdot\mathbb{I}[x\in S_{\text{old}}]}_{\text{Captured Reward}}+\underbrace{R(x)\cdot\mathbb{I}[x\notin S_{\text{old}}]}_{\text{Residual Reward}}$ (9)
As established in our analysis of Theorem 1, the "Captured Reward" term is precisely the expected estimate produced by the converged old GFlowNet, which we denote $\widehat{R}_{\text{old}}(x)\coloneqq\mathbb{E}[\widehat{R}_{\text{old}}(x;\tau^{\prime})]$ (cf. Eq. (8)), where the expectation is taken under the past GFlowNet’s backwards policy. This allows us to redefine the total reward as the sum of the old estimate and the residual target, $R_{\text{res}}(x)$:
$R(x)\approx\widehat{R}_{\text{old}}(x)+R_{\text{res}}(x).$
Enforcing the TB objective on the complement set is thus equivalent to enforcing a new balance condition over the entire set $\Gamma$, where the target is this $R_{\text{res}}(x)$.
This decomposition motivates two equivalent balance conditions for the new model, $\hat{R}_{\theta}(x;\tau)$:
@@
    Target-Residual (TR): The new model directly learns the residual target:
    $\hat{R}_{\theta}(x;\tau)=R(x)-\hat{R}_{\text{old}}(x).$ (10)
@@
    Flow-Additive (FA): The new model’s flow is added to the old flow to match the total reward (an algebraic rearrangement of the same principle):
    $\hat{R}_{\theta}(x;\tau)+\hat{R}_{\text{old}}(x)=R(x).$ (11)
While derived here for a single past model, this same principle applies when $\hat{R}_{\text{old}}(x)$ represents the combined flow of a pre-existing ensemble, where $\hat{R}_{\text{old}}(x)=\sum_{i}\hat{R}_{i}(x)$.
These two strategies are unified into a single Boosted Trajectory Balance Loss formulation, controlled by a hyperparameter $\alpha\in[0,1]$:
$\mathcal{L}_{\text{boost}}(\tau)=\left(\log\left[\frac{\hat{R}_{\theta}(x;\tau)+\alpha\hat{R}_{\text{old}}(x)}{R(x)-(1-\alpha)\hat{R}_{\text{old}}(x)}\right]\right)^{2}.$ (12)
@@
For a terminal $x$, let $\hat{R}_{\text{old}}(x)$ be an estimator of the target reward $R(x)$. Then:
@@
    (No Degradation) If $\hat{R}_{\text{old}}(x)=R(x)$, then the stationary points of $\mathcal{L}_{\text{boost}}$ (Eq. (12)) have $Z_{\theta}=0$.
@@
    (Residual Focus) If $\hat{R}_{\text{old}}(x)=0$, then the stationary points of $\mathcal{L}_{\text{boost}}$ are the same as those of the standard TB loss, $\mathcal{L}_{\text{TB}}$ (Eq. (4)).
@@
The TR condition, while theoretically sound, induces a undefined loss when $\hat{R}_{\text{old}}(x)\geq R(x)$, as the denominator in Eq. (12) becomes zero or negative. This problematic case occurs precisely when the old ensemble has perfectly learned (or over-estimated) a mode.
To mitigate this, we can derive a stable objective for this specific region. If we substitute the problematic $\hat{R}_{\text{old}}(x)$ with its converged target value, $R(x)$, directly into the general $\mathcal{L}_{\text{boost}}$ objective (Eq. (12)), the loss simplifies:
$\begin{split}\mathcal{L}_{\nabla}(\tau)&=\left(\log\left[\frac{\hat{R}_{\theta}(x;\tau)+\alpha R(x)}{R(x)-(1-\alpha)R(x)}\right]\right)^{2}\\ &=\left(\log\left[\frac{\hat{R}_{\theta}(x;\tau)}{\alpha R(x)}+1\right]\right)^{2}.\end{split}$ (13)
This simplified objective, which we refer to as $\mathcal{L}_{\nabla}(\tau)$ (cf. Eq. (13)), is well-defined and still correctly drives the new model’s flow $\hat{R}_{\theta}(x;\tau)$ to zero. We therefore use this as a practical safeguard for $\mathcal{L}_{\text{boost}}$ when its denominator is non-positive.
@@ 4.4 Sampling from the ensemble
The core principle of our framework is that each $i$-th stage is trained to capture a residual component of the total reward. Consequently, its learned partition function, $Z_{i}$, serves as an estimate of the total mass of that component. The natural way to sample from the full ensemble is therefore to treat it as a mixture model, where the probability of selecting the $i$-th stage is proportional to its learned mass, $Z_{i}$. The following theorem, proven in the appendix, formalizes that this sampling procedure correctly recovers the target distribution.
@@ 4.5 Relationship to Boosting VI
Notably, Boosted GFlowNets are also tightly connected to the framework of boosting variational inference (BVI; guo2016boosting; miller17boosting; locatello18boosting). To understand this relationship, we recall that BVI iteratively builds a mixture distribution by minimizing its Kullback-Leibler divergence to the target. That is, for each $t>1$,
$q^{\star},\beta^{\star}=\operatornamewithlimits{argmin}_{q\in\mathcal{H},\ \beta\in[0,1]}\mathcal{D}_{\mathrm{KL}}\left[(1-\beta)\cdot p^{(t)}+\beta\cdot q||\pi\right],$ (15)
for a tractable family $\mathcal{H}$ of distributions. Then, $p^{(t+1)}=(1-\beta^{\star})\cdot p^{(t)}+\beta^{\star}\cdot q^{\star}$. Clearly, each $p^{(t)}$ is a mixture of elements of $\mathcal{H}$ for $t>1$ (locatello2018boosting).
@@
Under the notations of Equation˜12, let $Z_{\mathrm{old}}$ and $Z_{\theta}$ be the normalizing constants for $\hat{R}_{\mathrm{old}}(x)\coloneqq\mathbb{E}[\hat{R}_{\mathrm{old}}(x;\tau^{\prime})]$ and $\hat{R}_{\theta}(x;\tau)$. Also, let $\beta=\nicefrac{{Z_{\theta}}}{{Z_{\mathrm{old}}+Z_{\theta}}}$ be the mixture weight in Equation˜15. Then, define $p_{\mathrm{old}}(\tau)\propto\hat{R}_{\mathrm{old}}(x)p_{B}(\tau|x)$ and $p_{\mathrm{tgt}}(\tau)=R(x)p_{B}(\tau|x)$ as the trajectory-level distributions induced by our current and target models, and let $p_{M}(\tau)=(1-\beta)\cdot p_{\mathrm{old}}(\tau)+\beta\cdot p_{F}(\tau)$ be the corresponding mixture distribution. In this scenario,
@@
Importantly, Proposition˜1 expands the connection between GFlowNets and VI (malkin2023gflownets) and provides a formal justification for the naming of our method.

### v2

# Boosted GFlowNets: Improving Exploration via Sequential Learning

@@ 1 Introduction
Generative Flow Networks (GFlowNets) learn stochastic policies over directed acyclic state graphs to sample structured objects in proportion to unnormalized terminal rewards (Bengio et al., 2021). This framework makes it possible to discover diverse candidates even when feedback is only available at the end of a trajectory, and has shown promise in scientific discovery and combinatorial design. In practice, however, training on large, multimodal targets is often bottlenecked by exploration: the learner rapidly concentrates on easy-to-reach modes while leaving hard-to-reach ones severely undercovered, limiting both coverage and generalization.
@@
Contributions. Our main contributions are:
  1. 1.
@@ 2 Related Work
GFlowNets face well-documented exploration bottlenecks on large, multimodal targets. Practical remedies span three main angles: (i) auxiliary exploration policies, e.g., sibling-augmented schemes that collect novelty-seeking trajectories, often via intrinsic rewards such as Random Network Distillation, and re-label them off-policy for training the main learner (Madan et al., 2025); (ii) adaptive curricula, where teacher–student mechanisms steer sampling toward under-covered regions identified by high loss or discrepancy signals (Kim et al., 2025); and (iii) loss-design perspectives, where alternative regression losses (zero-avoiding families) are motivated through divergence analyses to bias training away from mode collapse and toward broader coverage (Hu et al., 2025). Scaling-oriented approaches decompose the state graph and train subgraphs asynchronously to enlarge visited regions under fixed budgets (Silva et al., 2025b). Complementary theoretical works study conditions for correctness and distributional mismatch in finite-budget regimes, clarifying when coverage failures arise in practice (Silva et al., 2025a).
Boosting Variational Inference (BVI; Guo et al., 2016; Miller et al., 2017; Locatello et al., 2018b) generalizes classical boosting to probabilistic inference by iteratively constructing a mixture approximation to the target posterior. At each round, a new variational component is added to reduce the overall Kullback–Leibler divergence between the mixture and the true distribution. This procedure can be interpreted as performing functional gradient descent in the space of probability measures, where each component incrementally corrects the residual approximation error of the current mixture. Under mild smoothness and compactness assumptions on the variational family, BVI enjoys explicit convergence guarantees and systematically improves multimodal coverage by expanding the support of the approximation.
@@ 3 Background and Preliminaries
Generative Flow Networks (GFlowNets, Bengio et al. 2021) are probabilistic models designed to sample discrete objects $x\in\mathcal{X}$ with probability proportional to a non-negative reward function $R(x)$. Unlike standard generative models, which typically learn a normalized distribution, GFlowNets operate by training a pair of forward and backward policies over trajectories $\tau=(s_{0}\to s_{1}\to\dots\to s_{T})$ in a directed acyclic graph (DAG) of states. The forward policy $P_{F}$ induces a distribution over terminal states, while the backward policy $P_{B}$ provides a stochastic inverse, together defining a system of “flows” that approximate $R(x)$. This formulation has been shown effective in domains where rewards are sparse, multimodal, or expensive to evaluate, making GFlowNets a promising tool for scientific discovery and structured generative modeling.
Formally, let $\Gamma$ denote the set of all valid trajectories from the initial state $s_{0}$ to any terminal state $x\in\mathcal{X}$. Each trajectory $\tau\in\Gamma$ is assigned a forward probability $P_{F}(\tau)=\prod_{t=0}^{T-1}P_{F}(s_{t+1}\mid s_{t})$ and a corresponding backward probability $P_{B}(\tau\mid x)=\prod_{t=0}^{T-1}P_{B}(s_{t}\mid s_{t+1})$, where $s_{T}=x$. A GFlowNet is trained so that the marginal probability of sampling $x$ under the forward policy satisfies
@@
with $Z=\sum_{x\in\mathcal{X}}R(x)$ acting as the normalizing constant. The objective of training is therefore to align the induced terminal distribution of $P_{F}$ with the unnormalized target defined by $R(x)$.
Training objectives.
Several training objectives have been proposed to enforce the proportionality between $P_{F}(x)$ and $R(x)$, including Detailed Balance (DB, Malkin et al. 2022), Subtrajectory Balance (SubTB, Madan et al. 2023). Our work builds on the widely used Trajectory Balance condition (TB, Malkin et al. 2022). The TB objective introduces a scalar parameter $Z_{\theta}$ and enforces the following balance condition along full trajectories:
@@ 4.2 Induced Reward and Residual Principle
Let $\Gamma_{x}$ denote the set of trajectories that terminate at $x$. Throughout this section, we work with trained (hence fixed) GFlowNets and omit parameter dependence from the notation. For any GFlowNet $\mathfrak{g}:=(Z,P_{F},P_{B})$, terminal $x$, and trajectory $\tau\in\Gamma_{x}$, define the per-trajectory estimator
$\widehat{R}_{\mathfrak{g}}(x;\tau):=Z\,\frac{P_{F}(\tau)}{P_{B}(\tau\mid x)}.$ (6)
@@
$\widehat{R}_{\mathfrak{g}}(x):=\mathbb{E}_{\tau\sim P_{B}(\cdot\mid x)}\big[\widehat{R}_{\mathfrak{g}}(x;\tau)\big].$ (7)
@@
Fix a trained GFlowNet $\mathfrak{g}:=(Z,P_{F},P_{B})$, and define the support of its forward policy’s terminal distribution as
$S:=\{x\in\mathcal{X}:P_{F}(x)>0\}.$
Assume that the Trajectory Balance loss has zero expectation,
$\mathbb{E}_{\tau\sim P_{F}}\!\big[\mathcal{L}_{TB}(\tau)\big]=0.$
Further assume that, for every terminal $x\in S$, the backward policy $P_{B}(\cdot\mid x)$ and the restriction of the forward policy $P_{F}(\cdot)|_{\Gamma_{x}}$ are mutually absolutely continuous. Then, for every terminal $x\in\mathcal{X}$,
$\widehat{R}_{\mathfrak{g}}(x)=R(x)\,\mathbb{I}[x\in S]$
and
$\operatorname{Var}_{\tau\sim P_{B}(\cdot\mid x)}\!\big[\widehat{R}_{\mathfrak{g}}(x;\tau)\big]=0,$
@@
This yields the residual principle: given a frozen predictor that already matches parts of the target, its induced terminal estimate can be used to define a residual target. The next stage is then trained to capture this residual, thereby reallocating learning pressure toward under-covered regions while preserving coverage in regions that are already matched. This principle is operationalized in our boosted loss, detailed in the next subsection.
For an ensemble of trained GFlowNets $\mathcal{G}=\{\mathfrak{g}_{1},\dots,\mathfrak{g}_{N}\}$, we aggregate the induced estimates of its members via
$\widehat{R}_{\mathcal{G}}(x):=\sum_{n=1}^{N}\widehat{R}_{\mathfrak{g}_{n}}(x),$
To lighten notation, once the relevant trained model (single stage or ensemble) is clear from context, we omit the subscripts and simply write $\widehat{R}(x)$ and $\widehat{R}(x;\tau)$.
@@ 4.3 Boosted Trajectory Balance Loss
Our sequential training framework is motivated by a simple observation: the overall objective is to enforce the Trajectory Balance (TB) condition (Eq. (2)) over the set of all possible trajectories, $\Gamma$.
When a GFlowNet minimizes the expected Trajectory Balance (TB) loss, $\mathbb{E}_{\tau\sim P_{\mathrm{F}}}[\mathcal{L}_{\text{TB}}(\tau)]$ to zero the TB condition (Eq. (2)) is satisfied for all trajectories within the support of the forward policy, $P_{\mathrm{F}}$. Let $S$ be the set of terminal states covered by a converged past GFlowNet, and let $\Gamma_{S}\subset\Gamma$ be the set of trajectories ending in $S$. The remaining learning objective for a new GFlowNet is thus to enforce the same balance condition on the complement set of trajectories, $\Gamma\setminus\Gamma_{S}$, which lead to the under-covered modes.
@@
$R(x)=\underbrace{R(x)\cdot\mathbb{I}[x\in S]}_{\text{Captured Reward}}+\underbrace{R(x)\cdot\mathbb{I}[x\notin S]}_{\text{Residual Reward}}$ (8)
As established in our analysis of Theorem 1, the "Captured Reward" term is precisely the expected estimate produced by the trained GFlowNet, which we denote $\widehat{R}(x)$ (cf. Eq. (7)). This allows us to rewrite the equality above as:
$R(x)-\widehat{R}(x)=R(x)\mathbb{I}[x\notin S]$ (9)
Enforcing the TB condition on the complement is thus equivalent to enforcing a new balance condition over the entire set $\Gamma$, where the target is this $R(x)-\widehat{R}(x)$.
This decomposition suggests two equivalent balance conditions for the next stage, expressed in terms of the reward-induced estimator $\widehat{R}_{\theta}(x;\tau)\!\!\;:=Z_{\theta}\,\frac{P_{F}^{\theta}(\tau)}{P_{B}^{\theta}(\tau\mid x)}.$ The new model can be trained to satisfy, either:
@@
    Target-Residual (TR): the new stage directly matches the residual target,
    $\widehat{R}_{\theta}(x;\tau)=R(x)-\widehat{R}(x),$ (10)
@@
    Flow-Additive (FA): equivalently, its contribution adds to the frozen model to recover the full reward,
    $\widehat{R}_{\theta}(x;\tau)+\widehat{R}(x)=R(x),$ (11)
While derived here for a single past model, this same principle applies when $\widehat{R}$ represents the combined flow of a pre-existing ensemble.
These two strategies are unified into a single Boosted Trajectory Balance Loss, controlled by $\alpha\in[0,1]$ and the Monte Carlo budget $k$:
$\mathcal{L}_{\text{boost}}^{(k)}(\tau)=\left(\log\left[\frac{\widehat{R}_{\theta}(x;\tau)+\alpha\widehat{R}_{k}(x)}{R(x)-(1-\alpha)\widehat{R}_{k}(x)}\right]\right)^{2}.$ (12)
We omit the dependence on $k$ when it is clear from context or irrelevant to our analysis.
Here, $\widehat{R}_{k}(x)$ is a $k$-sample Monte Carlo estimate of the frozen model’s induced terminal estimate $\widehat{R}(x)$ (Eq. (7)). In the ensemble case, we draw $k$ trajectories from each member’s backward policy, average within each member, and sum the resulting estimates:
$\widehat{R}_{k}(x)\!:=\!\sum_{n=1}^{N}\frac{1}{k}\sum_{i=1}^{k}\widehat{R}_{\mathfrak{g}_{n}}(x;\tau_{n,i}),\tau_{n,i}{\sim}P_{B}^{(n)}(\cdot\mid x).$ (13)
@@
For a terminal $x$, let $\widehat{R}(x)$ be an estimator of the target reward $R(x)$. Then:
@@
    (No Degradation) If $\widehat{R}(x)=R(x)$, then the stationary points of $\mathcal{L}_{\text{boost}}$ (Eq. (12)) have $Z_{\theta}=0$.
@@
    (Residual Focus) If $\widehat{R}(x)=0$, then the stationary points of $\mathcal{L}_{\text{boost}}$ are the same as those of the standard TB loss, $\mathcal{L}_{\text{TB}}$ (Eq. (3)).
@@
The TR variant is mathematically undefined when $\widehat{R}(x)\geq R(x)$, as the denominator in Eq. (12) becomes zero or negative. This problematic case occurs precisely when the old ensemble has perfectly learned (or over-estimated) a mode. To mitigate this, we can choose $\alpha$ such that it remains as close to 0 while maintaining numerical stability.
@@ 4.4 Sampling from the ensemble
The core principle of our framework is that each stage $i$ is trained to capture a residual component of the total reward. Consequently, its learned partition function, $Z_{i}$, serves as an estimate of the total mass of that component. The natural way to sample from the full ensemble is therefore to treat it as a mixture model, where the probability of selecting stage $i$ is proportional to its learned mass, $Z_{i}$. The following theorem, proven in the appendix, formalizes that this sampling procedure correctly recovers the target distribution.
@@ 4.5 Relationship to Boosting VI
Notably, Boosted GFlowNets are also tightly connected to the framework of boosting variational inference (BVI; Guo et al., 2016; Miller et al., 2017; Locatello et al., 2018b). To understand this relationship, we recall that BVI iteratively builds a mixture distribution by minimizing its Kullback-Leibler divergence to the target. That is, for each $t>1$,
$q^{\star},\beta^{\star}=\operatorname*{arg\,min}_{q\in\mathcal{H},\ \beta\in[0,1]}\mathcal{D}_{\mathrm{KL}}\left[(1-\beta)\cdot p^{(t)}+\beta\cdot q||\pi\right],$ (15)
for a tractable family $\mathcal{H}$ of distributions. Then, $p^{(t+1)}=(1-\beta^{\star})\cdot p^{(t)}+\beta^{\star}\cdot q^{\star}$. Clearly, each $p^{(t)}$ is a mixture of elements of $\mathcal{H}$ for $t>1$ (Locatello et al., 2018a).
@@
Under the notations of Equation˜12, let $\widehat{Z}$ and $Z_{\theta}$ be the normalizing constants for $\widehat{R}(x)$ and $\hat{R}_{\theta}(x;\tau)$. Also, let $\beta=\nicefrac{{Z_{\theta}}}{{\widehat{Z}+Z_{\theta}}}$ be the mixture weight in Equation˜15. Then, define $\widehat{p}(\tau)\!\propto\!\widehat{R}(x)p_{B}(\tau|x)$ and $p_{\mathrm{tgt}}(\tau)\!=\!R(x)p_{B}(\tau|x)$ as the trajectory-level distributions induced by our current and target models, and let $p_{M}(\tau)=(1-\beta)\cdot\widehat{p}(\tau)+\beta\cdot p_{F}(\tau)$ be the corresponding mixture distribution. In this scenario,
@@
Importantly, Proposition˜1 expands the connection between GFlowNets and VI (Malkin et al., 2023) and provides a formal motivation for our method’s name.